[Hallc_running] [New Logentry] RC Daily update

bduran at jlab.org bduran at jlab.org
Wed Feb 19 11:55:01 EST 2020

Logentry Text:
Yesterday around 12:30 pm beam to the halls was resumed after the injector matching work. Following the beam checkout procedure we did an NMR measurement --> 47% longitudinal https://logbooks.jlab.org/entry/3786686 and started to take production data. We did NMR measurement every 5 hours. The polarization measurement results in order: 52.2% longitudinal  https://logbooks.jlab.org/entry/3786898 , 54.46% https://logbooks.jlab.org/entry/3787010, 55% https://logbooks.jlab.org/entry/3787175. 54.9% https://logbooks.jlab.org/entry/3787285

Today at 8:38am, we stopped taking data to start Moeller measurement which is ongoing right now. We'll resume production data taking.

Greg Smith is taking over RC after 1:30pm today.



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3787379
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200219/a6d0686a/attachment.html>

More information about the Hallc_running mailing list